OpenStudy (anonymous):

Restrict the function f(x)=x^2+4x+5 so that its inverse exists.

OpenStudy (anonymous):

make the graph

OpenStudy (anonymous):

ohhh haha thanks

OpenStudy (anonymous):

How would I find the inverse?

OpenStudy (anonymous):

define it on one side of the vertex only so first you have to find the vertex

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}=-\frac{4}{2}=-2\) so you could define your function on \(x>-2\) for example then to find the inverse, write \[x=y^2+4y+5\] and solve for \(y\)

OpenStudy (anonymous):

\[-2+\sqrt{x-1}\] ?